Proof of rational points on the general form of conic sections?


1
Expert's answer
2021-05-31T19:00:28-0400

 There is a theorem that states :

if a conic has 1 rational solution it has infinitely many rational solutions.


WE CAN PROVE IT AS FOLLOWS:


Suppose you have a rational conic as:

"ax^2+bxy+cy^2+dx+ey+f=0."


A generic line intersects a conic in 2 points.

If the conic and the line are rational {given by rational coefficients},


then the set consisting of those two points is invariant by Q-automorphisms,

so either each point is rational, either they are defined over a quadratic extension of Q

Q and the automorphism of that extension swaps the two intersection points.


If you have a rational point P on the conic,

then the rational lines going through it will intersect the rational conic at another rational point (possibly P itself if the line is tangent at P to the conic)

Moreover, two different lines going through P can only intersect at P


, so the lines that are not tangent to the conic at P

give you distinct points on the conic.


This actually gives you a bi-rational map between P1.

and the conic.


proof:


Suppose we have x,y "\epsilon" Q such that g(x, y) = 0. Consider the line through (x, y) with rational direction vector ( u ) parameterized by ( x,y )+t( u ). We claim that the second intersection point of the line and the conic is also a rational point (the first intersection point is (x, y), corresponding to t= 0.

g(x+tu,y+t) = a(x+tu)2+b(x+tu)(y+t)+c(y+t)2+

d(y+tu)+e(y+t) +fg(x,y) =0

= t2(au2+bu+c)+t(du+e+2aux+bx+byu+2y)


So the second intersection point corresponds (consider g(x+ tu, y+ t) = 0) to the rational parameter( t).


There are clearly infinitely many ways to choose u "\epsilon" Q such that t represents a nontrivial rational number, giving rise to infinitely many rational points on the curve of the form g(x+tu,y+t).